When a third number is averaged with 16 and 23, the result is 22. Find that third number.
Found 2 solutions by stanbon, wgomero:
Answer by stanbon(75887)

When a third number is averaged with 16 and 23, the result is 22. Find that third number.
------
Let the 3rd number be "x".
Then:
[16+23+x]/3=22
39+x=66
x=27 (The 3rd number is 27)
Cheers,
Stan H.

Answer by wgomero(28)

resolution:
let x be the unknown number
average of three number is (16 + 23 + x)/ 3 = 22; 16 + 23 + x = 66
x= 66 - 39; x = 27
the third number is 27.